# OTCliM: generating a near-surface climatology of optical turbulence strength ($C_{n}^{2}$) using gradient boosting

### v1

@@ Abstract
This study introduces OTCliM (Optical Turbulence Climatology using Machine learning), a novel approach for deriving comprehensive climatologies of atmospheric optical turbulence strength ($C_{n}^{2}$) using gradient boosting machines. OTCliM addresses the challenge of efficiently obtaining reliable site-specific $C_{n}^{2}$ climatologies, crucial for ground-based astronomy and free-space optical communication. Using gradient boosting machines and global reanalysis data, OTCliM extrapolates one year of measured $C_{n}^{2}$ into a multi-year time series. We assess OTCliM’s performance using $C_{n}^{2}$ data from 17 diverse stations in New York State, evaluating temporal extrapolation capabilities and geographical generalization. Our results demonstrate accurate predictions of four held-out years of $C_{n}^{2}$ across various sites, including complex urban environments, outperforming traditional analytical models. Non-urban models also show good geographical generalization compared to urban models, which captured non-general site-specific dependencies. A feature importance analysis confirms the physical consistency of the trained models. It also indicates the potential to uncover new insights into the physical processes governing $C_{n}^{2}$ from data. OTCliM’s ability to derive reliable $C_{n}^{2}$ climatologies from just one year of observations can potentially reduce resources required for future site surveys or enable studies for additional sites with the same resources.
@@ I Introduction
Atmospheric optical turbulence is highly relevant for optical ground-based astronomy and future free-space optical communication (FSOC). Both applications suffer from light getting distorted when propagating through the turbulent atmosphere. In astronomy, turbulent fluctuations of the atmospheric refractive index, known as optical turbulence (OT), cause blurry images and limit the detection of small objects [1]. FSOC links, which use optical beams to transmit data instead of traditional radio waves, experience reduced data rates or even link interruptions due to OT [2, 3]. Therefore, the optical turbulence strength ($C_{n}^{2}$) must be carefully considered in the design and operation of such optical systems. This requires robust statistical evaluation of $C_{n}^{2}$ over time for the operational sites of interest, known as a $C_{n}^{2}$ climatology. Such a climatology, derived from long-term $C_{n}^{2}$ data, portrays trends, seasonal variations, or potential anomalies in OT strength.
@@
However, conducting long-term site surveys at various locations of interest is time-consuming and resource-intensive. While $C_{n}^{2}$ can be obtained by post-processing numerical weather model output (e.g., Masciadri et al. [9]), running such models for multiple years is computationally expensive, and the accuracy of the resulting $C_{n}^{2}$ is very sensitive to the model configuration [10] and the selected $C_{n}^{2}$ parameterization [11]. We address these issues by proposing a novel machine learning-based (ML) approach called OTCliM (Optical Turbulence Climatology using Machine learning). OTCliM aims to extrapolate just one year of measured $C_{n}^{2}$ into a multi-year time series, enabling the generation of a comprehensive site-specific $C_{n}^{2}$ climatology with less data than a conventional site survey. As a result, site survey costs can potentially be reduced, climatologies can be obtained faster, and more sites can be surveyed within a given time frame. By leveraging ML, OTCliM can also model complex input-output relations, offering an advantage over traditional empirical $C_{n}^{2}$ models, which are discussed in the following section.
@@ II Relevant $C_{n}^{2}$ regression studies
Since the refractive index fluctuations quantified by $C_{n}^{2}$ are a result of density fluctuations due to temperature fluctuations, $C_{T}^{2}$ is linked to $C_{n}^{2}$ through the Gladstone equation
$C_{n}^{2}={\left(A\,P/\overline{T}^{2}\right)}^{2}C_{T}^{2}$ (2)
with $A\approx 7.9\times 10^{-5}\,\text{K}\,\text{hPa}^{-1}$ for optical wavelengths [16], station pressure $P$ in hPa, and mean air temperature $\overline{T}$ in K. Function $g(\zeta)$ in Eq. 1 is a similarity function, empirically determined from observations,
@@
The ML-based $C_{n}^{2}$ parameterizations in literature differ in their input variables, the ML regression technique employed, and if $C_{n}^{2}$ is measured at a single or multiple vertical levels. Much work has been devoted to single-level near-surface $C_{n}^{2}$ obtained as point measurements (e.g., from sonic anemometer) or via path-averaging (e.g., from scintillometer). Wang and Basu [18], for example, used fully connected neural networks to estimate single-level $C_{n}^{2}$ from routine meteorological data and gradients. Vorontsov et al. [19] also utilized neural networks but aimed at estimating path-averaged $C_{n}^{2}$ from images of a received laser beam that is broken up into speckles due to turbulence. Jellen et al. [20, 21] employed random forest and gradient boosting machines (GBM) to estimate path-averaged $C_{n}^{2}$ in a maritime environment from routine parameters and radiation fluxes. Most ML-based models do not yield analytical equations of the fitted model, unless they explicitly target it as shown by Bolbasova et al. [22]. The authors obtained a complex site-specific equation for single-level $C_{n}^{2}$, similar to non-ML empirical models proposed by, e.g., Sadot [23], Wang et al. [24], Arockia Bazil Raj et al. [25]. The difference between these empirical models and MOST-based models is that they are not derived physically but aim to capture the local optical turbulence strength well. Going beyond single-level $C_{n}^{2}$ estimates, Pierzyna et al. [26] proposed a physics-inspired framework to derive a data-driven similarity theory of optical turbulence that combines $C_{n}^{2}$ observations from multiple levels within the surface layer into a single non-dimensional model. In contrast to the near-surface studies, Su et al. [27] proposed a method to estimate a vertical $C_{n}^{2}$ profile by modeling OT strength as a function of radiosonde measurements. However, all these approaches require in-situ data, rendering them unsuitable for temporal extrapolation as targeted in the present study. Additionally, in-situ-based models need sensors deployed at the site of interest for $C_{n}^{2}$ estimation, which hinders their applicability to sites without instrumentation. To alleviate this issue, we proposed OTCliM, which parameterizes $C_{n}^{2}$ based on reanalysis data, which are available globally and for multiple decades into the past.
@@ III Methodology
Figure 1: Proposed OTCliM approach to extrapolate a measured 1-year time series of optical turbulence strength (yellow) to multiple years (orange) based on ERA5 reference data (blue). Robust yearly $C_{n}^{2}$ statistics can be obtained from the extrapolated data.
Our proposed OTCliM approach aims to extrapolate one year of observed $C_{n}^{2}$ to multiple years to obtain a robust statistical description of OT strength at a particular site. An overview of OTCliM is given in Fig. 1, which we base on the measure-correlate-predict (MCP) framework popular in wind energy [28, 29]. In the first step, $C_{n}^{2}$ is measured for one year (yellow) and then correlated to a reference dataset. For OTCliM, we utilize variables from the ERA5 reanalysis [30] extracted at the location of the $C_{n}^{2}$ measurements as reference data and gradient boosting machines (GBM) to regress the ERA5 time series to $C_{n}^{2}$ where they overlap (blue). In the second step, the trained model is utilized to extrapolate $C_{n}^{2}$ based on ERA5 to multiple years (prediction step in MCP), which can then be used to obtain site-specific seasonal statistics as presented in step 3.
@@ III.2 Baseline models
We employ two baseline models to put the performance of OTCliM into perspective. The first model is the traditional empirical W71 model as given by Eqns. 1 and 3. Since the similarity function $g(\zeta)$ does not contain site-specific model coefficients, the accuracy of the W71 $C_{n}^{2}$ estimates is expected to be lower than the site-specific OTCliM models. Secondly, we utilize GBM models trained with in-situ observations as input instead of reanalysis. Reanalysis datasets are coarse-resolution model outputs (spatial resolution typically $\mathcal{O}(10\,\text{km})$), which are expected to miss some local and complex processes and patterns modulating turbulence. Such processes would be contained in the in-situ data, so the performance of in-situ GBM models is expected to be higher than that of ERA5-based OTCliM models. These baselines allow us to disentangle the influence of the modeling approach (traditional vs. GBM) from that of the input data (ERA5 vs. in-situ). Appendix A gives more technical details about both baseline models.
@@ III.4 Performance evaluation
In Eq. 6, $\overline{\square}$ and $\sigma_{\square}$ are the mean and standard deviation of true ($\bm{y}$) and estimated ($\hat{\bm{y}}$) scaled $C_{n}^{2}$, respectively. Two validation strategies are employed to assess the temporal and geographical extrapolation capabilities of the OTCliM models. Note that only the testing data used to compute the scores differ between the evaluation strategies – the training data and the trained model remain the same.
@@
For the MCP application – the temporal extrapolation –, the model $\hat{f}_{s,t}$ trained on data from year $t\in\mathcal{T}$ needs to accurately predict OT strength at the same site for the held-out testing years $\mathcal{T}\setminus t$: $\hat{f}_{s,t}(\bm{x}_{i})=\hat{y}_{i}$ with $\bm{x}_{i}\in\bm{X}_{s,\mathcal{T}\setminus t}$ and $\hat{y}\in\bm{\hat{y}}_{s,\mathcal{T}\setminus t}$. The scores are computed between the true held-out data $\bm{y}_{s,\mathcal{T}\setminus t}$ and the predictions $\bm{\hat{y}}_{s,\mathcal{T}\setminus t}$ and are denoted $r_{s,t}$ and $\epsilon_{s,t}$. Training happens in a round-robin fashion, so one model is trained per training year and evaluated for the hold-out years. This process is repeated until all years are used for training once. The average performance of these models for a specific station $s$, called the MCP score, is obtained by averaging the scores across all training years: ¯$\overline{\square}_{s}={\langle\square_{s,t}\rangle}_{t\in\mathcal{T}}$.
@@ IV.1 New York State Mesonet (NYSM)
The New York State Mesonet (NYSM) comprises 127 standard weather stations spread across New York State, United States of America, and has been fully operational since 2018 [36]. These weather stations measure routine meteorological parameters such as 2 m-temperature, 10 m-wind speed and direction, surface pressure, and several other variables. The sampling rate of these measurements is on the order of seconds, and final values are reported as 10 min aggregates (mean and variance). However, measuring $C_{n}^{2}$ requires higher-frequency observations to resolve the inertial range of turbulence. Such high-frequency measurements are available at a subset of 17 stations – the flux stations – additionally equipped with Campbell Scientific CSAT3A sonic anemometers mounted at 9 m height. The NYSM sonic anemometers measure the three wind components and the sonic temperature at $f_{s}=10$ Hz [36], which is high enough to obtain $C_{n}^{2}$.
@@
Structure function approach
As noted earlier, $C_{n}^{2}$ quantifies the strength of refractive index fluctuations due to density fluctuations caused by turbulent temperature fluctuations. Therefore, $C_{n}^{2}$ can be obtained by measuring the temperature fluctuations quantified by the coefficient of the second-order temperature structure function $C_{T}^{2}$, which is defined in the inertial range as
@@
Quality assurance and control
Since our GBM models are purely data-driven, the quality of the fitted model strongly depends on the quality of the training data, so we apply several QA/QC steps to obtain a clean dataset. First, $C_{T}^{2}$ has to be determined from the slope of the inertial range, i.e., slopes of $2/3$. Since the inertial range is not always resolved or well captured in every 5 min window, only $C_{T}^{2}$ values with well-fitted slopes are kept ($2/3\pm 5\%$ slope, $R^{2}>0.95$). A similar procedure was followed by He and Basu [38] for the analysis of simulated data. Second, $S_{T}^{2}$ in Eq. 8 is defined spatially as a function of separation distance $\Delta x$, but the sonic anemometer measures $T$ temporally. Taylor’s frozen turbulence assumption is commonly made to convert the temporal sonic signal into a spatial form using the horizontal mean wind speed $\overline{M}$ and $\Delta x=\overline{M}/f_{s}$. However, Taylor’s assumption might break down in low wind conditions, so all $C_{T}^{2}$ values for $\overline{M}<1$ m s-1 are discarded. Finally, precipitation strongly affects sonic temperature measurements [39], so only $C_{T}^{2}$ values for dry conditions can be used. This filter is also justified from an operational perspective because optical telescopes are typically not operated in the rain, and FSOC links are strongly attenuated by snow [40]. While the QA/QC procedure discards $\sim 50$% of the data throughout the year, it still contains ca. fifty thousand samples per site and training year spread throughout diverse meteorological conditions, which we view as adequate.
Distribution scaling

### v2

@@ Abstract
This study introduces OTCliM (Optical Turbulence Climatology using Machine learning), a novel approach for deriving comprehensive climatologies of atmospheric optical turbulence strength ($C_{n}^{2}$) using gradient boosting machines. OTCliM addresses the challenge of efficiently obtaining reliable site-specific $C_{n}^{2}$ climatologies near the surface, crucial for ground-based astronomy and free-space optical communication. Using gradient boosting machines and global reanalysis data, OTCliM extrapolates one year of measured $C_{n}^{2}$ into a multi-year time series. We assess OTCliM’s performance using $C_{n}^{2}$ data from 17 diverse stations in New York State, evaluating temporal extrapolation capabilities and geographical generalization. Our results demonstrate accurate predictions of four held-out years of $C_{n}^{2}$ across various sites, including complex urban environments, outperforming traditional analytical models. Non-urban models also show good geographical generalization compared to urban models, which capture non-general site-specific dependencies. A feature importance analysis confirms the physical consistency of the trained models. It also indicates the potential to uncover new insights into the physical processes governing $C_{n}^{2}$ from data. OTCliM’s ability to derive reliable $C_{n}^{2}$ climatologies from just one year of observations can potentially reduce resources required for future site surveys or enable studies for additional sites with the same resources.
@@ I Introduction
Atmospheric optical turbulence is highly relevant for optical ground-based astronomy and future free-space optical communication (FSOC). Both applications suffer from light getting distorted when propagating through the turbulent atmosphere. In astronomy, turbulent fluctuations of the atmospheric refractive index, known as optical turbulence (OT), cause blurry images and limit the detection of small objects [1]. FSOC links, which use optical beams to transmit data instead of traditional radio waves, experience reduced data rates or even link interruptions due to OT [2, 3]. Therefore, the optical turbulence strength ($C_{n}^{2}$, where index $n$ denotes the refractive index) must be carefully considered in the design and operation of such optical systems. This requires robust statistical evaluation of $C_{n}^{2}$ over time for the operational sites of interest, known as a $C_{n}^{2}$ climatology. Such a climatology, derived from long-term $C_{n}^{2}$ data, portrays trends, seasonal variations, or potential anomalies in OT strength.
@@
However, conducting long-term site surveys at various locations of interest is time-consuming and resource-intensive. While $C_{n}^{2}$ can be obtained by post-processing numerical weather model output (e.g., Masciadri et al. [9]), running such models for multiple years is computationally expensive, and the accuracy of the resulting $C_{n}^{2}$ is very sensitive to the model configuration [10] and the selected $C_{n}^{2}$ parameterization [11]. We address these issues by proposing a novel machine learning-based (ML) approach called OTCliM (Optical Turbulence Climatology using Machine learning). OTCliM aims to extrapolate just one year of measured near-surface $C_{n}^{2}$ into a multi-year time series, enabling the generation of a comprehensive site-specific $C_{n}^{2}$ climatology with less data than a conventional site survey. Our approach does not yet provide full vertical $C_{n}^{2}$ profiles as often sought in astronomy or FSOC. But for near-surface conditions, which are highly relevant for, e.g., horizontal near-surface FSOC links, site survey costs can potentially be reduced, climatologies can be obtained faster, and more sites can be surveyed within a given time frame. By leveraging ML, OTCliM can also model complex input-output relations, offering an advantage over traditional empirical $C_{n}^{2}$ models, which are discussed in the following section.
@@ II Relevant $C_{n}^{2}$ regression studies
Sensible heat flux $\overline{w^{\prime}\theta^{\prime}}$ and friction velocity $u_{*}={\left(\overline{u^{\prime}w^{\prime}}^{2}+\overline{v^{\prime}w^{\prime}}^{2}\right)}^{1/4}$ describe the vertical (i.e., surface-normal) turbulent transport of heat ($\theta^{\prime}$) and momentum ($u^{\prime}$, $v^{\prime}$) due to fluctuations of the vertical wind component ($w^{\prime}$) [16]. As these parameters capture the two effects modulating turbulence – buoyancy (sensible heat flux) and wind shear (friction velocity) –, they are commonly used in turbulence parameterizations. Since refractive index fluctuations ($C_{n}^{2}$) are driven by density fluctuations due to temperature and moisture fluctuations, $C_{n}^{2}$ is linked to $C_{T}^{2}$ as [17, 18]
$C_{n}^{2}={\left(A\,\frac{P}{\overline{T}^{2}}\right)}^{2}{\left(1+\frac{0.03}{\beta}\right)}^{2}\,C_{T}^{2}$ (2)
with $A\approx 7.9\times 10^{-5}\,\text{K}\,\text{hPa}^{-1}$ for optical wavelengths [19], station pressure $P$ in hPa, and mean air temperature $\overline{T}$ in K. The contribution of moisture fluctuations to $C_{n}^{2}$ is captured by the Bowen ratio $\beta$, the ratio of the sensible heat flux and the latent heat flux due to evaporation. If $C_{n}^{2}$ is estimated above the boundary layer in the free atmosphere, a modified formulation of Eq. 2 can be used [20].
Function $g(\zeta)$ in Eq. 1 is a similarity function, empirically determined from observations,
@@
Figure 1: Proposed OTCliM approach to extrapolate a measured 1-year time series of optical turbulence strength (golden yellow) to multiple years (orange) based on ERA5 reference data (blue). Robust yearly $C_{n}^{2}$ statistics can be obtained from the extrapolated data.
The ML-based $C_{n}^{2}$ parameterizations in literature differ in their input variables, the ML regression technique employed, and if $C_{n}^{2}$ is measured at a single or multiple vertical levels. Much work has been devoted to single-level near-surface $C_{n}^{2}$ obtained as point measurements (e.g., from sonic anemometer) or via path-averaging (e.g., from scintillometer). Wang and Basu [22], for example, used fully connected neural networks to estimate single-level $C_{n}^{2}$ from routine meteorological data and gradients. Vorontsov et al. [23] also utilized neural networks but aimed at estimating path-averaged $C_{n}^{2}$ from images of a received laser beam that is broken up into speckles due to turbulence. Jellen et al. [24, 25] employed random forest and gradient boosting machines (GBM) to estimate path-averaged $C_{n}^{2}$ in a maritime environment from routine parameters and radiation fluxes. Most ML-based models do not yield analytical equations of the fitted model, unless they explicitly target it as shown by Bolbasova et al. [26]. The authors obtained a complex site-specific equation for single-level $C_{n}^{2}$, similar to non-ML empirical models proposed by, e.g., Sadot [27], Wang et al. [28], or Arockia Bazil Raj et al. [29]. The difference between these empirical models and MOST-based models is that they are not derived physically but aim to capture the local optical turbulence strength well. Going beyond single-level $C_{n}^{2}$ estimates, Pierzyna et al. [30] proposed a physics-inspired framework to derive a data-driven similarity theory of optical turbulence that combines $C_{n}^{2}$ observations from multiple levels within the surface layer into a single non-dimensional model. In contrast to the near-surface studies, Su et al. [31] proposed a method to estimate a vertical $C_{n}^{2}$ profile by modeling OT strength as a function of radiosonde measurements. Cherubini et al. [32] presented an ML framework to estimate seeing, which is the vertical integral of the $C_{n}^{2}$ profile, close to the surface and in the free atmosphere, but does not explicitly resolve vertical levels. Milli et al. [33] tackled yet another estimation angle by using ML to forecast future OT conditions for the next two hours based on past observations and auxiliary features. However, all these approaches require in-situ data, rendering them unsuitable for temporal extrapolation as targeted in the present study. Additionally, in-situ-based models need sensors deployed at the site of interest for $C_{n}^{2}$ estimation, which hinders their applicability to sites without instrumentation. To alleviate this issue, we proposed OTCliM, which parameterizes $C_{n}^{2}$ based on reanalysis data, which are available globally and for multiple decades into the past.
@@ III Methodology
Our proposed OTCliM [^1] The Python code for training is available on GitHub: https://github.com/mpierzyna/otclim approach aims to extrapolate one year of observed $C_{n}^{2}$ to multiple years to obtain a robust statistical description of OT strength at a particular site. An overview of OTCliM is given in Fig. 1, which we base on the measure-correlate-predict (MCP) framework popular in wind energy [35, 36]. In the first step, $C_{n}^{2}$ is measured for one year (golden yellow) and then correlated to a reference dataset. For OTCliM, we utilize variables from the ERA5 reanalysis [37] extracted at the location of the $C_{n}^{2}$ measurements as reference data and gradient boosting machines (GBM) to regress the ERA5 time series to $C_{n}^{2}$ where they overlap (blue). In the second step, the trained model is utilized to extrapolate $C_{n}^{2}$ based on ERA5 to multiple years (prediction step in MCP), which can then be used to obtain site-specific seasonal statistics as presented in step 3.
@@ III.2 Baseline models
We employ two baseline models to put the performance of OTCliM into perspective. The first model is the traditional empirical W71 model as given by Eqns. 1 and 3. Since the similarity function $g(\zeta)$ does not contain site-specific model coefficients, the accuracy of the W71 $C_{n}^{2}$ estimates is expected to be lower than the site-specific OTCliM models. Secondly, we utilize GBM models trained with in-situ observations as input instead of reanalysis. Reanalysis datasets are coarse-resolution model outputs (spatial resolution typically $\mathcal{O}(10\,\text{km})$), which are expected to miss some local and complex processes and patterns modulating turbulence. Such processes would be contained in the in-situ data, so the performance of in-situ GBM models is expected to be higher than that of ERA5-based OTCliM models. These baselines allow us to disentangle the influence of the modeling approach (traditional vs. GBM) from that of the input data (ERA5 vs. in-situ). Appendix C gives more technical details about both baseline models.
@@ III.4 Performance evaluation
In Eq. 6, the overbar denotes the mean and $\sigma$ the standard deviation of true ($\bm{y}$) and estimated ($\hat{\bm{y}}$) scaled $C_{n}^{2}$, respectively. Two validation strategies are employed to assess the temporal and geographical extrapolation capabilities of the OTCliM models. Note that only the testing data used to compute the scores differ between the evaluation strategies – the training data and the trained model remain the same.
@@
For the MCP application – the temporal extrapolation –, the model $\hat{f}_{s,t}$ trained on data from year $t\in\mathcal{T}$ needs to accurately predict OT strength at the same site for the held-out testing years $\mathcal{T}\setminus t$: $\hat{f}_{s,t}(\bm{x}_{i})=\hat{y}_{i}$ with $\bm{x}_{i}\in\bm{X}_{s,\mathcal{T}\setminus t}$ and $\hat{y}\in\bm{\hat{y}}_{s,\mathcal{T}\setminus t}$. The scores are computed between the true held-out data $\bm{y}_{s,\mathcal{T}\setminus t}$ and the predictions $\bm{\hat{y}}_{s,\mathcal{T}\setminus t}$ and are denoted $\square_{s,t}$, where $\square$ is a placeholder for the performance metrics $r$ and $\epsilon$. Training happens in a round-robin fashion, so one model is trained per training year and evaluated for the hold-out years. This process is repeated until all years are used for training once. The average performance of these models for a specific station $s$, called the MCP score, is obtained by averaging the scores across all training years: ¯$\overline{\square}_{s}={\langle\square_{s,t}\rangle}_{t\in\mathcal{T}}$.
@@ IV.1 New York State Mesonet (NYSM)
The New York State Mesonet (NYSM) comprises 127 standard weather stations as of 2024 spread across New York State, United States of America, and has been fully operational since 2018 [43]. These weather stations measure routine meteorological parameters such as 2 m-temperature, 10 m-wind speed and direction, surface pressure, and several other variables. The sampling rate of these measurements is on the order of seconds, and final values are reported as 10 min aggregates (mean and variance). However, measuring $C_{n}^{2}$ requires higher-frequency observations to resolve the inertial range of turbulence. Such high-frequency measurements are available at a subset of 17 stations – the flux stations – additionally equipped with Campbell Scientific CSAT3A sonic anemometers mounted at 9 m height. The NYSM sonic anemometers measure the three wind components and the sonic temperature at $f_{s}=10$ Hz [43], which is high enough to obtain $C_{n}^{2}$.
@@
IV.1.1 Structure function approach
As noted in the context of Eq. 2, $C_{n}^{2}$ quantifies the strength of refractive index fluctuations due to density fluctuations caused by turbulent temperature and moisture fluctuations. The temperature $T_{s}$ measured by sonic anemometers also contains a humidity contribution as $T_{s}=T(1+0.51q)$ with specific humidity $q$ [45]. We assume that using the simplified version
$C_{n}^{2}\approx{\left(A\,P/\overline{T}^{2}\right)}^{2}C_{T_{s}}^{2}$ (8)
of Eq. 2 implicitly accounts for moisture if the strength of these sonic temperature fluctuations, $C_{T_{s}}^{2}$, is used. That is because $C_{T_{s}}^{2}$ can be decomposed as $C_{T_{s}}^{2}\approx C_{T}^{2}+1.02\overline{T}C_{Tq}+0.26\overline{T}^{2}C_{q}^{2}$ with $C_{T}^{2}$ and $C_{q}^{2}$ representing the strength of pure temperature and moisture fluctuations and $C_{Tq}$ the cross-term between the two. For readability, we will refer to $C_{T_{s}}^{2}$ simply as $C_{T}^{2}$ but would like to reiterate that a moisture contribution is included implicitly.
In the inertial range, $C_{T}^{2}$ is the coefficient of the second-order structure function of the (sonic) temperature defined as
@@ IV.1.1 Structure function approach
IV.1.2 Quality assurance and control (QA/QC)
Since our GBM models are purely data-driven, the quality of the fitted model strongly depends on the quality of the training data, so we apply several QA/QC steps to obtain a clean dataset. First, $C_{T}^{2}$ has to be determined from the slope of the inertial range, i.e., slopes of $2/3$. Since the inertial range is not always resolved or well captured in every 5 min window, only $C_{T}^{2}$ values with well-fitted slopes are kept ($2/3\pm 5\%$ slope, $R^{2}>0.95$). A similar procedure was followed by He and Basu [46] for the analysis of simulated data. Second, $S_{T}^{2}$ in Eq. 9 is defined spatially as a function of separation distance $\Delta x$, but the sonic anemometer measures $T$ temporally. Taylor’s frozen turbulence assumption is commonly made to convert the temporal sonic signal into a spatial form using the horizontal mean wind speed $\overline{M}$ and $\Delta x=\overline{M}/f_{s}$. However, Taylor’s assumption might break down in low wind conditions, so all $C_{T}^{2}$ values for $\overline{M}<1$ m s-1 are discarded. Finally, precipitation strongly affects sonic temperature measurements [47], so only $C_{T}^{2}$ values for dry conditions can be used. This filter is also justified from an operational perspective because optical telescopes are typically not operated in the rain, and FSOC links are strongly attenuated by snow [48].
Being an observed dataset, the sonic temperature signal contains gaps due to, e.g., power or communication outages or instrument malfunctioning [49], leading to $\sim 16$% of missing values on average. That leaves $\sim 84$% of the 5-min non-overlapping windows gap-free and suitable for the computation of $C_{T}^{2}$. The three QA/QC steps further reduce the average number of valid samples, leaving $\sim 66\%$ after inertial range fitting, $\sim 55\%$ after ensuring the applicability of the Taylor assumption, and, $\sim 53\%$ after discarding precipitation events. Although only half of the whole dataset is considered valid, it still contains ca. fifty thousand samples per site and training year spread throughout diverse meteorological conditions. This claim is supported in Appendix A, where we present that the QA/QC procedure generally maintains the ratio between stable and unstable atmospheric conditions before and after QA/QC. Thus, the final dataset is considered adequate for training the OTCliM models.
IV.1.3 Distribution scaling